A researcher studying the nutritional value of a new candy places a 4.60-gram sample of the candy inside a bomb calorimeter and combusts it in excess oxygen. The observerd temperature increase is 2.76°C. If the heat capacity of the calorimeter is 48.60 kJ·K-1 , how many nutritional Calories are there per gram of the candy?
1
Expert's answer
2017-10-23T05:38:19-0400
dT(1g) = 2.76/4.6 = 0.6 °C (0.6 °K)
E = 48.6kJ*K^-1 * 0.6 °K = 29.16kJ